A box having a mass of 80.0kg is dragged across a rough horizontal floor by means of a rope tied on the front of it. The coefficient of friction between the box and the floor is 0.450. If the angle between the rope and the floor is 38.0 degrees, what force must be exerted on the rope to move the box at a constant velocity ? 


1
Expert's answer
2021-05-05T10:53:50-0400

"\vec{F}+m\vec{g}+\vec{N}+\vec{F_{fr}}=\vec{0},"

"F-mg\sin\alpha-\mu mg\cos\alpha=0,"

"F=mg(\sin\alpha+\mu\cos\alpha)=760~N."